OpenStudy (anonymous):

Write a recursive formula for the sequence 15, 26, 48, 92, 180, …. Then find the next term.

OpenStudy (mathstudent55):

26 - 15 = 1 * 11 48 - 26 = 22 = 2 * 11 92 - 48 = 44 = 4 * 11 180 - 92 = 88 = 8 * 11

OpenStudy (mertsj):

\[a _{n}=2a _{n-1}-4\]

OpenStudy (mathstudent55):

\(a_1 = 15\) \(a_2 = a_1 + 11(2 - 1)\) \(a_3 = a_2 + 11(3 - 1) \) \(a_n = a_{n - 1} + 11(n - 1)\)

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (anonymous):

\[a _{n}=11*2^n+4,n \in I,n \ge 0\]